OpenStudy (anonymous):

Joanne bought 3 baseball cards for $160. If the first card was as expensive as the combined price of the second and the third cards, and the second card cost $20 more than the third card, what is the price of the first card? A)80 B)50 C)100 D)60

OpenStudy (anonymous):

I believe it might be 80 or 100 but I am not sure

hero (hero):

f = price of 1st card s = price of 2nd card t = price of 3rd card f + s + t = 160 f = s + t s = t + 20

OpenStudy (anonymous):

Then it would be 80 correct? 80+50+30=160 80=50+30 50=30+20

hero (hero):

Correct